Prove that the relation R:{(x,y)|x-y is divisible by 3} is an Equivalence relation.(R is



defined over Z) L2 CO2 10

1
Expert's answer
2022-01-03T16:27:06-0500

Let us prove that the relation "R=\{(x,y)|\ x-y\text{ is divisible by 3}\}\subset\Z\times\Z" is an equivalence relation.

Since "x-x=0" is divisible by 3 for any "x\in \Z," we conclude that "(x,x)\in R" for any "x\in \Z," and hence the relation is reflexive.

Let "(x,y)\in R." Then "x-y" is divisible by 3. It follows that "y-x=-(x-y)" is also divisible by 3, and thus the relation "R" is symmetric.

Let "(x,y),(y,z)\in R." Then "x-y" is divisible by 3 and "y-z" is divisible by 3. It follows that "x-z=(x-y)+(y-z)" is also divisible by 3, and therefore "(x,z)\in R." We conclude that "R" is a transitive relation.

Therefore, the relation "R=\{(x,y)|\ x-y\text{ is divisible by 3}\}\subset\Z\times\Z" is an equivalence relation.